### v1

# Constructing Benchmarks and Interventions for Combating Hallucinations in LLMs

Abstract
Large language models (LLMs) are susceptible to hallucination, which sparked a widespread effort to detect and prevent them. Recent work attempts to mitigate hallucinations by intervening in the model’s computation during generation, using different setups and heuristics. Those works lack separation between different hallucination causes. In this work, we first introduce an approach for constructing datasets based on the model knowledge for detection and intervention methods in closed-book and open-book question-answering settings. We then characterize the effect of different choices for intervention, such as the intervened components (MLPs, attention block, residual stream, and specific heads), and how often and how strongly to intervene. We find that intervention success varies depending on the component, with some components being detrimental to language modeling capabilities. Finally, we find that interventions can benefit from pre-hallucination steering direction instead of post-hallucination. The code is available here.
1 Introduction
A notable issue identified in LLMs is their susceptibility to generate outputs that are not grounded in the input of the model or in real-world facts, or are inconsistent with earlier generations in a session (Ji et al., 2023; Sharma et al., 2023; Kalai & Vempala, 2023). These different issues are typically titled collectively as hallucinations. The importance of addressing these phenomena lies in the reduced reliability of models that sometimes hallucinate. Numerous efforts have been made to tackle these issues, including blackbox approaches that propose suprervised and reinforcement learning finetuning recipes attempting to fix the generated text (Tian et al., 2023; Razumovskaia et al., 2023; Ouyang et al., 2022), as well as whitebox approaches, which aim to detect and prevent hallucinations using the model’s inner states, such as weights and internal activations (Burns et al., 2022; He et al., 2023; Rateike et al., 2023; Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023).
In this work we focus on whitebox approaches, which may enable earlier actions upon hallucination during response generation and propose a better understanding of the hallucination mechanism within LLMs. We claim that a proper benchmark for the task of whitebox intervention should only include hallucinations for cases in which the model “knows” the correct response but still generates it incorrectly, and not cases in which, for example, the model was not exposed to the ground truth knowledge at pretraining, and therefore has no chance to respond correctly. To this end, we first define the different hallucination types w.r.t effective whitebox mitigation. We then propose a method to automatically differentiate between these types and with it construct model-specific benchmarks (each benchmark is tailored to a given model’s internal knowledge). We construct such benchmarks in two settings, one that requires the model to rely on knowledge from the input context within the prompt (termed open-book hallucination) and another that requires the model to rely only on its parametric knowledge (termed closed-book hallucination).
The scenario of closed-book hallucination, wherein the model hallucinates despite knowing the correct answer and lacking any alternative answer in the context, is highly intriguing. Intuitively, one might assume that such hallucinations should not occur. However, examples in the literature demonstrate how user input can trigger such behavior (see Nardo (2023); Wolf et al. (2023)). In Section 3.3, we propose a novel setting to investigate this phenomenon further.
Armed with tailored benchmarks, we turn to whitebox analysis and utilize steering vectors (TurnTrout et al., 2023; Subramani et al., 2022) to mitigate hallucinations. Different from prior work, which detected and created a steering vector for hallucinations after they were generated (Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024), we target detection before hallucination onset, with the goal of steering the generation away from hallucination before it is materialized in surface tokens. In addition, despite some success in detection and mitigation in prior work, there are no clear guidelines on how to do it. We use our framework to create such guidelines by rigorously investigating different choices for intervention such as on which component in the model’s architecture to apply them. We introduce dynamic intervention, performing it only when needed, and compare it to static intervention. As well as using pre-hallucination inner activations for the steering vector instead of post-hallucination. We discover several insights: (i) classification accuracy is not a good approximation to generation accuracy and perplexity should be measured as well; (ii) intervening on the attention component is preferred over the MLP part; (iii) intervention on the residual has a perplexity compromise, which is mitigated to some extent when intervening dynamically;(iv) intervening using pre-hallucination inner states for the steering vector is better than post-hallucination;
Our main contributions are: (I) A methodology for constructing per-LLM benchmarks containing grounded and hallucinated outputs in cases where the model knows the information for a correct response in its parameters, in both open-book and closed-book settings. We will release the datasets for our tested models (Llama2-7b and Goat-7B); (II) A comprehensive investigation into optimal intervention strategies within the model architecture. and a novel dynamic intervention approach that can help intervene only when needed. (III) Insights into the evaluation of interventions.
2 Problem Setup
Consider a question-answering (QA) setup, where a model is given a question and generates an answer. We consider two settings: open-book QA where the gold-truth answer appears in a given context and closed-book QA where the gold-truth answer resides in the model’s parameters. A common recipe for mitigating hallucinations involves the following steps: (1) constructing a dataset with hallucination labels; (2) training a detector (classifier) on intermediate representations to identify cases of hallucinations; and (3) mitigation - intervening in the model’s computation to prevent a hallucinated answer. The third step requires the detector from step two to decide where to intervene. In this work we focus on steps (1) and (3). See Appendix D for details on how we perform step (2), training the detector.
2.1 Mitigation Setup
We aim to mitigate hallucinations by adjusting the model’s activations. Following prior work in this area (Li et al., 2023; Hoscilowicz et al., 2024; Chen et al., 2024; Marks & Tegmark, 2023; Zou et al., 2023) and in other tasks (Bau et al., 2019; Ravfogel et al., 2021), our intervention approach is to add a vector to the model’s activations , at some layer and component , where . For the MLP, attention block, and residual, our modification is at the output of the layer, and for the heads, it is at the output of the layer before applying the projection matrix back to the embedding space. The vector is meant to define a linear separation between hallucination and grounded subspaces. Then, during inference, our intervention formula, , shifts the activation vector in the direction of the grounded subspace, hopefully making the model generate a grounded answer. Here emphasizes the impact of the intervention. is sometimes known as a steering vector (Subramani et al., 2022; TurnTrout et al., 2023). We note that steering vectors do not introduce supplementary knowledge to the model, and computing them typically relies on labeled data. This information will be used in deciding how to construct the dataset.
Our paper explores the gory details of the mitigation step, how to compute the steering vector, evaluate its performance, and suggest insights for efficient mitigation.
2.2 Knowledge Categorization
In closed-book setting Hallucinations can be caused by different reasons, depending on whether or not the model knows the golden response, i.e., the model is expected to induce the golden response tokens from its activations with high confidence. We define three types of knowledge categories that in each hallucination can occur:
1.
The model does not know the golden answer at all.
2.
The model considers the golden answer as one of a few possible ones, without a clear preference for the golden answer.
3.
The model knows the golden answer.
We argue that if the model does not know the golden answer (type 1), it is not possible to prevent hallucination without using external knowledge. If the model has not clear preference for the golden answer (type 2), intervention without external knowledge might be possible, but seems unlikely. Instead, the model could choose to refrain from answering in cases of uncertainty in the answer. This is a different type of intervention that we defer for future work. On the other hand, if the model does know the golden answer, it may be possible to direct its output towards the golden answer via whitebox intervention. Therefore, we focus on type 3 in this work. The automatic way we differentiate between those types of knowledge is explained in Section 3.
This categorization into hallucination from the model knowledge that we propose was not done in prior work on mitigation (Zhang et al., 2024; Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024), and therefore prior benchmarks mixed the 3 types.111One exception is a concurrent work byYuan et al. (2024), which tries to consider the model’s knowledge in creating a dataset. We advocate for the significance of this categorization to address the diverse causes of hallucination effectively.
3 Intervention Dataset Construction
Training and evaluating detectors and steering vectors for mitigating hallucinations necessitates labeled datasets, containing hallucinated and grounded examples. We specifically aim at constructing such datasets for type 3 hallucinations (Section 2.2). To this end, we first describe an automatic method to differentiate between the three hallucination types. Then, we utilize this method to create training, validation, and test sets in open-book and closed-book QA settings for type 3 cases. Figure 1 outlines this process.
3.1 Hallucination Example Selection
To automatically differentiate between the three types of knowledge situations that can create hallucinations, we use repeated sampling from a tested LLM . Concretely, given a QA example consisting of a question and a gold answer, we sample answers from using temperature sampling, and categorize the example according to the number of times the gold response was sampled: 0 times refers to type-1, 1–3 times refers to type-2, and 4–5 times refers to type-3. We retain only type-3 examples, thus tailoring a dataset per model . Here and onwards in the paper, we will use the terminology for type-3 examples as the model knows the answer. From those examples, we will next create a labeled hallucination dataset. See Appendix A.1 for specifics.
This method is easy to implement and we found it useful for splitting into our three types.222One may also train a dedicated model to separate examples by knowledge of a model (Kadavath et al., 2022) , but that is more complicated than our approach. We applied it in both open- and closed-book settings; see below.333This is primarily needed in the closed-book setting but to maintain consistency we applied it in both settings.
3.2 Open-Book Dataset Creation
In this scenario, we study hallucinations that occur when there is a conflict between a model’s parametric knowledge and a given context. A hallucination in this case is when the model’s answer deviates from information given in the prompt context. We use the DisentQA dataset Neeman et al. (2022), where each example is a question , a text context , and two answers: a parametric answer and a contextual answer . See Figure 2 for an example.
question: how many chromosomes does a human diploid cell have? context: Humans are diploid organisms , carrying two complete sets of chromosomes : one set of 23 chromosomes from their father and one set of 23 chromosomes from their mother . The two sets combined provide a full complement of 2 chromosomes . This total number of chromosomes is called the chromosome number . The zygotic number is defined as the number of chromosomes in zygotic cells . Human zygotes are diploid , hence with a zygotic number of 2 . answer: Parametric answer (): 46. Contextual answer (): 2
We keep only examples for which the model knows the parametric answer, meaning it was generated at least 4 out of 5 times when sampling given only the question without the additional context.
Next, given a full DisentQA example {}, containing a question and context inputs, and the parametric and contextual gold answers, we measure how much the model prefers over : . Where represents the rank assigned by the model to token compared to all other tokens after projection to the vocabulary space at the last layer; a low-rank number indicates a higher preference for answer . If for some threshold then the example is labeled ‘grounded’, and if it is labeled ‘hallucination’. The result of this process is a model-specific open-book DisentQA labeled for hallucinations. See full statistics and more details in Appendices A.1, A.2 and B.
3.3 Closed-Book Dataset Creation
In this setting, we study hallucinations that occur when an LLM generates an answer that deviates from the world knowledge it has in its parameters.444Note that type 2 could be problematic to use as the model does not clearly prefer the golden answer thus the steering vector may not steer to the golden answer. To this end, we follow prior work that shows that models may simulate different personas, making them hallucinate even though they know the answer (Wolf et al., 2023; Joshi et al., 2023; Janus, 2022). This kind of behavior may happen when a user writes something untrue or misleading (Nardo, 2023; Wolf et al., 2023; Zhang et al., 2023; Xu et al., 2023).
Concretely, we constructed 20 bad-shot examples using ChatGPT (Ouyang et al., 2022). These are false QA pairs, where the false answer is semantically similar to the gold one. For instance, here is a good-shot example and its corresponding bad-shot example:
- Good-shot:
question: Which element has the chemical symbol ’H’? answer: Hydrogen
- Bad-shot:
question: Which element has the chemical symbol ’H’? wrong answer: Helium
Three good-shots (or three bad-shots) question: Which colour lies between red and yellow in the visible spectrum? (wrong) answer: Golden answer (): Orange
Then, for a given LLM , we used to answer the questions in TriviaQA (Joshi et al., 2017) in a closed-book setup. As part of the question instruction, we provide the good-shots or bad-shots as few-shot exemplars for the task (Figure 3). For a given example {, } of a question and a gold answer , we label it as ‘grounded’ or ‘hallucination’ via the difference in rank of following the bad-shot or good-shot prompts: . If the rank does not change ( ), the example is considered grounded. If the worsens ( ) and the model predicted the answer correctly given the good-shot, we consider the example as a hallucination. Otherwise, the example is removed. This labeling procedure is illustrated in Figure 4. More details, as well as examples of generated hallucinations using bad shots, are in Appendices A.1, A.3 and B.
4 Intervention Analysis Approach
With our datasets labeled with hallucinations of type 3, where the model knows the answer but still makes a mistake, we perform a rigorous analysis of whitebox model intervention using steering vectors (See Section 2.1). To this end, we discuss the design choices of intervention via steering vectors in this setting, and reassess proper evaluation metrics, indicating that both classification and generation accuracy are important to measure.
4.1 Exploring Methodological Variability in Intervention
Various studies employed different settings for intervention, and it remains unclear which approaches are most effective. We conduct experiments with the common choices that appear in most works to establish guidelines, as well as introduce novel settings. Figure 5 presents the variables we explore. Specifically, we identify the following degrees of freedom from the body of work in designing an intervention experiment.
First, a steering vector may intervene in different components: MLP (Zhang et al., 2024), attention block (Zhang et al., 2024), heads (Chen et al., 2024; Li et al., 2023; Hoscilowicz et al., 2024), or residuals (Marks & Tegmark, 2023; Zou et al., 2023)555Zou et al. (2023) also used the attention for a different intervention using LoRRA.. In this work, we will intervene in all the components.
Second, the steering vector is commonly computed using examples containing the generated answer (Li et al., 2023; Hoscilowicz et al., 2024; Chen et al., 2024; Marks & Tegmark, 2023; Zhang et al., 2024; He et al., 2023; Zou et al., 2023). However, it is then used for mitigation before the answer is generated. Thus, there may be a discrepancy between the way it was computed and how it is used. In our experiments, we consider both the common post-answer method and a new pre-answer setup where the steering vector is calculated before answer generation instead of after. Pre-answer is a novel setting we introduce, which is more aligned with how is used during intervention (before hallucination is generated).
Third, the intervention (-compute) may be computed from a detection classifier or from another summary statistic of hallucination vs. grounded cases. In particular, it is common to take the linear separator for the former (Marks & Tegmark, 2023; Chen et al., 2024; Li et al., 2023) and the difference in mean vectors of grounded and hallucination examples for the latter (Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024).666One exception is Zhang et al. (2024), who trained an auto-encoder to learn . In both cases, the decision of where to intervene is made using the detector classifier.
In addition, we introduce another degree of freedom. In static intervention occurs in the same components and layers, for all examples, usually by picking the places that got the highest detection score. This is a common strategy (Li et al., 2023; Marks & Tegmark, 2023; Zhang et al., 2024; He et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024). However, such drastic intervention may not always be necessary. We suggest a different option: dynamic intervention, in which the components to intervene are chosen in each example differently. Our intuition is that intervention should be applied when hallucination is likely to occur, which is pointed at by our detector. Thus, given a threshold , we intervene only in components whose detector accuracy is above (a high-quality detector) and the detector classified the example as hallucination for them.
We also explored the difference in intervention effect between fine-tuned and pre-trained models and investigated both open-book and closed-book settings.
4.2 Evaluating Intervention Impact
Recent work lacks a consistent evaluation method for measuring intervention success. Most compare probabilities of the golden answer and other possible answers (Marks & Tegmark, 2023; Zou et al., 2023; Li et al., 2023; Zhang et al., 2024; Chen et al., 2024; Hoscilowicz et al., 2024), while others also generate answers and evaluate them with a language model (Li et al., 2023; Zhang et al., 2024; Chen et al., 2024).
We advocate for examining multiple metrics to provide a holistic picture of intervention effects. To measure intervention success, we report the classification accuracy. In the open-book case, we measure how often the model prefers the contextual answer over the parametric answer by comparing Rank()Rank(). In the closed-book case, we compare the rank of the golden answer given a bad-shot prompt and a good-shot prompt (Rank bad shotsRank good shots). See Appendix C for detailed information.
The classification accuracy is simple and intuitive. However, recent work suggests that comparing probabilities of answer tokens is not a good proxy for text generation (Wang et al., 2024). We argue that this discrepancy is particularly crucial for intervention assessment. Therefore, we report the generation accuracy, which examines whether the desired answer was generated in greedy decoding of 20 tokens. Note, that we focus on a generation of a few tokens, as the model may produce a lengthy correct answer that should also be classified as grounded.
Finally, to ensure that the intervention has not compromised the model’s performance, we report its perplexity on 100 random articles from WikiText-103 (Merity et al., 2016) before and after an intervention. We call this Wiki ppl. See additional details in Appendix C.777We also tried pure perplexity but it was less stable than Wiki ppl.
5 Implementation Details
In all experiments, we randomly selected 1000 examples from each label (hallucinate/grounded) for analysis in each dataset, and split to 70%/10%/20% for training/validation/test.888Result on the general distribution of the data show similar but weaker trends; see Appendix F.4. We use a linear classifier (SVM) for detection, similar to prior work (Li et al., 2023; Marks & Tegmark, 2023). All results are averaged across three experiments, each with a different random seed, random split into training, validation, and testing, and the classifier seed in detection experiments. Detection results and the generalization of the detection in different settings can be found in Appendices D and F.5.
We evaluate Llama-2-7B (Touvron et al., 2023) in most experiments. We also evaluate GOAT-7B (Bekbayev et al., 2023), a fine-tuned version of Llama-2, in a subset of the experiments, to highlight differences between base and fine-tuned models.
To explore the various options raised in Section 4.1, we start with mass-mean direction intervention with dynamically and pre-answer input to , a configuration that showed promising results in initial exploration. We then modify a single variable at a time (dynamic/static intervention, computing the steering vector pre/post the answer etc.) to assess its impact on performance in an ablation setup. Additional hyper-parameter details are in Appendix E.999When comparing interventions with mass-mean vs. classifier direction, we found inconsistent results with no clear pattern. See discussion in Appendix F.3..
6 Intervention Results
Figures 6(a) and 7(a) show the accuracy on the open-book DisentQA and closed-book TriviaQA settings w.r.t intervening with increasing values of , which controls the magnitude of the intervention. We see different behaviors in the two settings. While in closed-book classification and generation accuracy are coupled, in open-book, when is increased, classification accuracy mostly increases but generation peaks and then drops, indicating that classification accuracy is not always a good approximation for generation accuracy, which is what we care for when using the model language abilities. We also see drop in Wikipedia perplexity as grows (Figures 6(c) and 7(c)), indicating worse language modeling capability with stronger intervention. This occurs in both settings but with different changes per setting. Therefore, evaluating also perplexity is recommended when assessing intervention performance.
Component-wise, we see that closed-book and open-book also behave differently. In open-book, attention and heads are best performing and very similar, under classification, generation and perplexity evaluation. In closed-book, residual has better classification and generation accuracy, but it declines at higher and perplexity levels are increasing quickly too, while attention accuracy keeps improving without any perplexity drop.
In Section 4.1 we proposed a novel pre-answer intervention strategy. Figure 8 presents a comparison between pre-answer and post-answer intervetion strategies101010We run this test only on the open-book setting as it has a clear answer ( for grounded examples and for hallucinate examples).. We can see that while on classification accuracy, the two strategies are on par, under generation and perplexity, it is clear that pre-answer is a better strategy. In addition, under post-answer, intervention at the heads component is best performing, but under pre-answer, attention component gains and matches heads performance.
We conclude that a fine intervening component in both settings is the attention, and to reach good results intervention should be performed using the pre-answer strategy.
6.1 Static Interventions vs. Dynamic Intervention
This section compares static and dynamic interventions. As discussed in Section 4.1, static intervention is the common approach in the literature and modifies the same components in every input example. This intervention usually selects the places with the highest detection to always intervene in. Dynamic intervention is a new approach we propose, which decides which components to modify given an input example based on the detector’s decision. Dynamic intervention might be preferable since unnecessary interventions might do more harm than good (the steering vector is designed to move the representation between subspaces, using it is not designed for cases where the representation is already in the "correct" subspace).
As Table 1 shows, in both settings, there is minimal difference between the two approaches in most cases with a slight overall preference for the static intervention. However, when intervening in the residual, dynamic intervention yields significantly better results, while static intervention leads to poor generation accuracy and perplexity. Again, we see that classification accuracy is not correlated with generation accuracy, this time in both settings.
To further illustrate the effect on the residual component, we provide examples of generated text using static vs. dynamic interventions that demonstrate how static intervention in the residual component negatively impacts generation in Appendix F.1.
We conclude that when intervening in the residual it is preferred to use the dynamic intervention and elsewhere it varies with generally slight preference for the static intervention.
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
| Attention | static | ||||||
| dynamic | |||||||
| MLP | static | ||||||
| dynamic | |||||||
| Heads | static | ||||||
| dynamic | |||||||
| Residual | static | ||||||
| dynamic | |||||||
| None | 0 | ||||||
6.2 Finetuning Facilitates Intervention
In this section, we explore the impact of fine-tuning on intervention success by comparing two models: Llama2-7B vs. Goat-7B, a fine-tuned version of Llama2.111111We decided to leave chat models for future work as they require modifying the prompt to a user/assistant format instead of a regular prompt, making direct comparisons of base and chat models difficult. Table 2 shows the average change resulting from intervention compared to no intervention. (Full results are found in Appendix F.2.) Notably, the intervention significantly enhances the generation and classification accuracy of Goat compared to Llama2 across most components, without negatively affecting perplexity. This suggests the advantage of utilizing the fine-tuned model over its basic counterpart, as shown in Li et al. (2023) but does not correspond to other works (Zhang et al., 2024). More research is needed to characterize differences in intervention effects on pre-trained and fine-tuned models.
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
| Attention | Llama2 | ||||||
| Goat | |||||||
| MLP | Llama2 | ||||||
| Goat | |||||||
| Heads | Llama2 | ||||||
| Goat | |||||||
| Residual | Llama2 | ||||||
| Goat | |||||||
7 Related Work
Prior work has considered cases of LM failures under various terms such as factuality (Chuang et al., 2023; He et al., 2023), lying (Azaria & Mitchell, 2023; Pacchiardi et al., 2023), truthfulness (Marks & Tegmark, 2023; Li et al., 2023; Zhang et al., 2024; Zou et al., 2023), sycophancy (Sharma et al., 2023), and as hallucinations (Shi et al., 2023; Dhuliawala et al., 2023). Hallucinations are sometimes defined as cases of model mistakes that seem plausible to a user (Ji et al., 2023).
An important distinction is between extrinsic (closed-book) and intrinsic (open-book) hallucinations (Ji et al., 2023; CH-Wang et al., 2023). The interventionist literature has mostly studied the closed-book case, while we consider both cases.
Research on hallucination typically falls into two main categories: detection and mitigation. Detecting hallucinations can involve treating the model as a black box, and posing questions or sampling (Gekhman et al., 2023; Pacchiardi et al., 2023; Manakul et al., 2023). Another line of work has tried to detect hallucinations or factuality by examining the model’s hidden representations, often by training a detection classifier (Burns et al., 2022; He et al., 2023; Rateike et al., 2023; Slobodkin et al., 2023; Azaria & Mitchell, 2023; CH-Wang et al., 2023; Yuksekgonul et al., 2023; Chen et al., 2023; Yin et al., 2024) We similarly use detectors in this work, with an emphasis on detection before a hallucination occurs.
Mitigating hallucinations can be done via prompting (Si et al., 2022; Dhuliawala et al., 2023; Lei et al., 2023) or fine-tuning (Tian et al., 2023; Razumovskaia et al., 2023). Other methods involve modifying the model’s logits (Chuang et al., 2023; Kalai & Vempala, 2023). Our work focuses on addressing hallucinations by intervening in the inner layers of the model (Zhang et al., 2024; Li et al., 2023; Marks & Tegmark, 2023; Zou et al., 2023; Chen et al., 2024; Hoscilowicz et al., 2024), which potentially does not require external resources or longer prompts. In this work, we characterize the effect of different choices for intervention, both from prior work and novel techniques we propose.
8 Discussion and Conclusion
In this work, we established a framework for assessing white-box hallucination mitigation techniques in open-book and closed-book settings. We propose a typology of hallucination types from the lens of a model’s knowledge, and highlight the feasibility of type-3 hallucinations – cases where the model knows the answer but does not generate it. We discussed the significance of these aspects and developed a methodology for creating appropriate datasets with hallucination labels.
We performed a comprehensive investigation of different variables involved in interventionist approaches for mitigating hallucinations. Notable insights include the importance of computing intervention vectors before reading an answer (input to ); the difference between evaluating classification and generation accuracy and the importance of perplexity evaluation; the potential of dynamic interventions that vary by example with its main importance in residual intervention; and the pros and cons traits of different components – while intervening in the residual reduces hallucinations, it also compromises the model’s language modeling capabilities. In comparison, intervention in the attention component consistently performs well across various measures and datasets.
This study establishes a framework for evaluating white-box hallucination mitigation techniques across open-book and closed-book settings, including a typology of hallucination types based on the model’s knowledge, while uncovering insights into effective intervention strategies.
9 Limitations
This work considers a wide range of possible configurations when mitigating hallucinations via white-box interventions. However, one limitation is that we only focused on one model (Llama2-7B) and one of its fine-tuned variants (Goat-7B). Additionally, we presented results on one open-book dataset (DisentQA) and one closed-book dataset (TriviaQA), both in question-answering setups. Lastly, we use an initial setting and modify a single variable each time instead of checking across all options of modifications. To form a broader understanding it is essential to investigate additional settings and models.
Appendix A Dataset Construction Specifics
A.1 General Dataset Construction Specifics
We created at least 1,000 examples labeled as “hallucinate" and 1,000 as “grounded". We filtered out instances where the golden/parametric/contextual answers exceeded 5 tokens or where the prompt exceeded 1,000 tokens.
To assess the model’s knowledge of the golden answer, we employed a beam search generation of 20 tokens, repeated 5 times with a temperature of 0.8. We utilized only examples falling under type 3 (Section 3.1), ensuring each answer was generated at least four times.
A.2 Open-Book Dataset Construction Specifics
We loaded the DisentQA (open-book) dataset from https://docs.google.com/document/d/1Z4vA7ifMQTk5YBF3BEYCFSnIvXCYaznLP_7VBcXPEeU/edit#heading=h.3prdi119z3tn and used <P> type counterfactual examples from the dataset with the name v10-simplified_simplified-nq-train_factual_counterfactual_disentangled_baseline_train_split.csv.
In the open book setting, to determine the model’s rank in generating the parametric/contextual answer, we computed the difference in rank of the first token that differs in the answer tokens between the two answers. And used 1 as the threshold .
A.3 Closed-Book Dataset Construction Specifics
In the closed-book setting, the rank was determined using the average rank of all answer tokens. Additionally, given the extensive size of TriviaQA, we randomly sampled 20,000 examples to construct our labeled dataset and loaded it from the "datasets" library. Table 3 shows examples of generated hallucinations of Llama2 using bad-shots in greedy decoding.
| Prompt | Grounded | Hallucinated |
[…]
Lastly, the full list of the 20 bad shots and good shots:
Good shots:
question: What is the capital of France?
answer: Paris
question: How many continents are there?
answer: 7
question: Who wrote ’Romeo and Juliet’?
answer: William Shakespeare
question: What is the square root of 64?
answer: 8
question: Which element has the chemical symbol ’H’?
answer: Hydrogen
question: Who was the first President of the United States?
answer: George Washington
question: What is the powerhouse of the cell?
answer: Mitochondria
question: In what year did World War II end?
answer: 1945
question: What is the currency of Japan?
answer: Japanese Yen
question: Who painted the Mona Lisa?
answer: Leonardo da Vinci
question: What is the speed of light?
answer: 299,792 kilometers per second
question: How many sides does a hexagon have?
answer: 6
question: What is the boiling point of water in Celsius?
answer: 100 degrees
question: Who wrote ’To Kill a Mockingbird’?
answer: Harper Lee
question: What is the capital of Australia?
answer: Canberra
question: What is the largest ocean on Earth?
answer: Pacific Ocean
question: Who discovered penicillin?
answer: Alexander Fleming
question: What is the chemical symbol for gold?
answer: Au
question: What is the smallest prime number?
answer: 2
question: How many planets are there in our solar system?
answer: 8
Bad shots:
question: What is the capital of France?
answer: Berlin
question: How many continents are there?
answer: 6
question: Who wrote ’Romeo and Juliet’?
answer: Jane Austen
question: What is the square root of 64?
answer: 7
question: Which element has the chemical symbol ’H’?
answer: Helium
question: Who was the first President of the United States?
answer: Abraham Lincoln
question: What is the powerhouse of the cell?
answer: Golgi Apparatus
question: In what year did World War II end?
answer: 1939
question: What is the currency of Japan?
answer: Euro
question: Who painted the Mona Lisa?
answer: Pablo Picasso
question: What is the speed of light?
answer: 300,000 kilometers per second
question: How many sides does a hexagon have?
answer: 5
question: What is the boiling point of water in Celsius?
answer: 50 degrees
question: Who wrote ’To Kill a Mockingbird’?
answer: J.K. Rowling
question: What is the capital of Australia?
answer: Sydney
question: What is the largest ocean on Earth?
answer: Atlantic Ocean
question: Who discovered penicillin?
answer: Isaac Newton
question: What is the chemical symbol for gold?
answer: Ag
question: What is the smallest prime number?
answer: 1
question: How many planets are there in our solar system?
answer: 9
Appendix B Final Datasets Statistic
In section 3 we created a labeled dataset given a model and a dataset. DisentQA for the open-book dataset and TriviaQA for the closed-book dataset. The statistic of the final dataset count of the number of examples in each label can be found in Tabel 4, the else label is all the examples that did not classify to the other two labels. We can see that the different settings have different results, indicating the importance of creating a labeled dataset regarding the specific model.
| Data name | # Grounded | # hallucinate | # else |
| closed book-Llama2-TriviaQA | 4553 | 1451 | 10306 |
| open book-Llama2-DisentQA | 1673 | 2869 | 8958 |
| closed book-Goat-TriviaQA | 2618 | 1507 | 12185 |
| open book-Goat-DisentQA | 3019 | 1261 | 9220 |
Appendix C Specific Evaluation Measures
In the open book setting, we assess classification accuracy by comparing the model’s preference for the , Rank()Rank() Regarding the closed book setting, we compare Rank( | bad shots) Rank( | good shots). In both settings, the rank calculation mirrors that of the dataset creation step. If the equations are true the example is labeled ‘grounded’, else ‘hallucinate’.
Generation accuracy in the open book setting evaluates whether the was generated and the was not. In the closed book setting evaluates whether the golden answer is produced.
Lastly, to calculate the Wiki ppl we use 100 random articles and from each, we use the first 100 tokens to calculate the perplexity.
Appendix D Detection Results
Figures 9(a) and 9(b) show hallucination detection results on the (open-book) DisentQA-based dataset and (closed-book) TriviaQA-based dataset, respectively, when applied to different model components at different layers, using Llama-2-7B.
On most layers, detection is significantly higher than random (dashed lines), indicating a signal of possible hallucination prior to the occurrence of hallucination. Additionally, we observe that the different components yield fairly similar results, with a slight advantage for detection based on the residual stream.
We remind the reader that detecting possible hallucinations before they occur is a more challenging task than detecting them post-hoc (CH-Wang et al., 2023). Indeed, if we concatenate the answer to the prompt, we reach over 90% accuracy in most layers and components, see Figure 13. Given that in practical use cases the answer is not known ahead of generation, we emphasize the importance of pre-hoc detection.
For completeness, we present the detection results on one data split of the validation set on the residual, attention, and MLP layers in Figures 10 and 11, utilizing three different classifiers: a linear classifier trained on the vectors as they are (labeled classifier in the figures), a linear classifier trained only on the vectors’ norms (labeled norm in the figures), and a linear classifier trained on the normalized vectors (labeled direction in the figures). It’s evident that the linear classifier trained on the normalized vectors (direction classifier) achieves the best results across most components, layers, and datasets.
Additionally in Figure 12 we see the classification results on the heads on Llama2-7B in the two settings. We can see that those results are a bit less good than the ones on the other model components with much better results in the open-book setting than the closed-book setting.
Lastly, in Figures 14 and 15, we observe the direction classifier’s accuracy across the Goat-7B model’s layers on both datasets. The detection results are similar and slightly better than those on Llama2.
Appendix E Hyperparams for Intervention
We aim to determine the optimal hyperparameters through validation and subsequently apply these parameters to the test set. We will select based on the average score on the classification accuracy on the validation set (200 total) using one seed. We optimized for the classification accuracy since this is the criterion according to which the data is labeled.
For the dynamic version, we will test the TD threshold parameter (TD = 0.65 / 0.7), while for the static version, we will explore the number of layers and heads to intervene on (5 layers and 50 heads / 3 heads and 30 heads).
The best hyperparameters for Llama2 on both DisentQA and TriviaQA are TD = 0.65 for the dynamic intervention and 5 layers and 50 heads for the static version. Full results are provided in Tables 5, 6, 7 and 8.
We also found that for the Goat model, the best TD is 0.65 for the dynamic intervention on both datasets. The full results are presented in Tables 9 and 10.
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 60.5 | 59.0 |
| MLP | 56.5 | 54.0 |
| Heads | 60.0 | 51.0 |
| Residual | 62.5 | 61.0 |
| None | 50.0 | 50.0 |
| Average | 57.9 | 55.0 |
| component | score (3 layers) | score (5 layers) |
| Attention | 53.0 | 63.0 |
| MLP | 49.5 | 52.0 |
| Heads | 50.0 | 51.5 |
| Residual | 83.0 | 83.5 |
| None | 50.0 | 50.0 |
| Average | 57.1 | 60.0 |
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 55.5 | 50.0 |
| MLP | 48.5 | 50.0 |
| Heads | 51.0 | 50.0 |
| Residual | 66.5 | 55.5 |
| None | 50.0 | 50.0 |
| Average | 54.3 | 51.1 |
| component | score (3 layers) | score (5 layers) |
| Attention | 58.0 | 56.0 |
| MLP | 46.5 | 48.5 |
| Heads | 49.5 | 50.0 |
| Residual | 68.5 | 82.0 |
| None | 50.0 | 50.0 |
| Average | 54.5 | 57.3 |
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 63.5 | 62.0 |
| MLP | 62.0 | 61.5 |
| Heads | 65.0 | 61.5 |
| Residual | 62.0 | 64.5 |
| None | 50.0 | 50.0 |
| Average | 60.5 | 59.9 |
| component | score (TD=0.65) | score (TD=0.7) |
| Attention | 54.0 | 49.5 |
| MLP | 57.0 | 50.5 |
| Heads | 48.5 | 49.5 |
| Residual | 70.5 | 50.0 |
| None | 50.0 | 50.0 |
| Average | 56.0 | 49.9 |
Appendix F Additional Results
In this section, we show a few additional results to the ones in Section 6.
F.1 Generation Examples of Static and Dynamic Intervention
To complete the picture from Section 6.1 we show a few more examples of intervention on the residual using dynamic intervention vs. static intervention on Llama2-7B in Table 11 and 12. We can see that the impact of the static intervention has made the text much less comprehensive. The impact on the DisentQA dataset is much stronger than on the TriviaQA dataset, where even the dynamic intervention creates a less comprehensive text.
| Prompt | Dynamic generation | Static generation |
[…]
| Prompt | Dynamic generation | Static generation |
[…]
F.2 Additional Pre-training vs. Finetuing Results
In Section 6.2, we demonstrated the improvement resulting from the intervention to evaluate the disparity between Llama2 and Goat. To provide a comprehensive view, we present in Table 13 the complete results instead of solely the improvement from the baseline. It is evident from the data that Goat achieves superior results following the intervention.
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
| Attention | Llama2 | ||||||
| Goat | |||||||
| MLP | Llama2 | ||||||
| Goat | |||||||
| Heads | Llama2 | ||||||
| Goat | |||||||
| Residual | Llama2 | ||||||
| Goat | |||||||
| None | Llama2 | ||||||
| Goat | |||||||
F.3 Mass-mean Direction vs. Classifier Direction
We aim to explore the distinction between a mass-mean direction and a classifier direction. The classifier direction option employs the direction generated by the linear classifier. To ensure that the two methods differ only in direction, we normalized the classifier steering vectors to the mass-mean norm. The results are presented in Table 14. Overall, both options yield similar results, with a slight overall advantage observed for the classifier direction.
| Setting | DisentQA | TriviaQA | |||||
| component | direction type | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity | Classification Accuracy | Generation Accuracy | Wikipedia Perplexity |
| Attention | mass-mean | ||||||
| classifier | |||||||
| MLP | mass-mean | ||||||
| classifier | |||||||
| Heads | mass-mean | ||||||
| classifier | |||||||
| Residual | mass-mean | ||||||
| classifier | |||||||
| None | 0 | ||||||
F.4 Results on Weighted Dataset
Here are the results of dynamic intervention with on a weighted-labeled dataset. This dataset comprises the test set consisting of 200 hallucinate examples, 200 grounded examples, and 200 other examples, weighted by the count of each label in the full dataset. This approach aids in estimating how intervention will impact a generally distributed dataset. The results are depicted in Figures 16 and 17. These findings align with the results obtained on a dataset evenly split between hallucination and grounded, as presented in the main paper, albeit with weaker trends.
F.5 Generalization of Detection
So far we have performed each experiment within a given dataset and model. Do hallucination detectors generalize across datasets or models? In Figure 18(a), we employ a classifier trained on Goat-7B and evaluate its performance when detecting hallucinations from Llama2-7B states, on the DisentQA dataset (GoatLlama2; DisentQA). Interestingly, the Goat classifier consistently outperforms random chance across all layers, suggesting a notable correlation between the inner states of the two models. These results correspond to related works (Wu et al., 2020). However, the results are much lower than a detector trained and evaluated on Llama2 states (Llama2Llama2; 9(a)), indicating that cross-model generalization is limited.
In contrast, Figure 18(b) illustrates that detectors struggle to generalize across the two datasets. This underscores that to help understand and mitigate hallucinations it is important to address each hallucination scenario individually and make cautious claims about each one.
Similar to these findings, we observe a comparable outcome on the Llama2 TriviaQA setting in Figure 19. Additionally, the results of using the Llama2 model to learn a classifier for the Goat model, which is similar to the results in the other direction, also show accuracy above random, are in Figure 20. Lastly, in Figure 21, we present the results on the Goat model using each dataset to detect hallucinations on the other dataset.